OpenStudy (anonymous):

whats the answer for 8/3x=-4/15

OpenStudy (anonymous):

8/3x=-4/15 8/12x=-1/15 2/3x=-1/15 -30/3=x -10=x x=-10

OpenStudy (anonymous):

thanks what about this one4/5 - 1/2x=x+5/6x

OpenStudy (anonymous):

Follow the procedure thjones demonstrated to solve for x.